Write the standard form of the equation Parabola with vertex at (-2,4) and focus at (3,4)
Answer by stanbon(75887)

Write the standard form of the equation Parabola with vertex at (-2,4) and focus at (3,4)
----
Plot those 2 points and you will see that the parabola opens
to the right.
----
Form: (y-k)^2 = 4p(x-h)
---
p = 3--2 = 5; so 4p = 20
---
h = -2 and k = 4
-----
Equation:
(y-4)^2 = 20(x+2)
=====================
Cheers,
Stan H.
======================